Given a level of confidence of 95% and a population standard deviation of 9, answer the following:
(A) What other information is necessary to find the sample size (n)?
(B) Find the Maxium Error of Estimate (E) if n=72. Show all work.
=============================================================================
Here is what I stated doing. Please correct if I am wrong.
a: n= sqrt=9=3
Please Help me!!

Answer by stanbon(75887)

Given a level of confidence of 95% and a population standard deviation of 9, answer the following:
(A) What other information is necessary to find the sample size (n)?
n = [z*s/E]^2
You need to know the maximum allowed error.
-----------------------------------
(B) Find the Maximum Error of Estimate (E) if n=72. Show all work.
ME = z*s/sqrt(n) = 1.96*9/sqrt(72) = 2.08
=====================
Cheers,
Stan H.